Timing distribution
Slow Controls infrastructure and GUI
Vertical Slice preparedness
Online algorithm benchmark

Priorities with Lab Access:

Characterize Timing Stability (10-DTC-daisy chain or star-fanout-topology)
Debug DTC Event Building
Horizontal Slice Testing (detector subsystem priority is Vertical Slice Testing)
Retransmission testing

Rick -- Almost done with event headers. Then will work on backpressure feature during event building between DTCs.. Eric says IPMI is fixed (was on NOvA network).

Eric -- FCC3 was warm when Eric was there (74F). Making progress with Giani and Sophie. Sophie and Eric had ART module to Visualizer working.

Antonio -- Completed first draft of EPICS-otsdaq interface document. Back to work on usability of Web Dashboard.

Bertrand -- pushed online algo changes to clustering. Clustering is down to 0.3ms (1000x optimized over last few years). First piece also simplified which may make it possible in the FPGA. Bottlenecks now are data products and movement. Maybe at 4ms for single thread. Operations review went well - no recommendations for TDAQ Ops.

Glenn -- DCS Daq15 RAID is configured now. Posgres accounts setup. Readonly and Readwrite account for EPICS db. Daq15 still at FCC.

Lei -- CRV are still thinking about event fixed page size.

Ralf -- Proposed changes to OFFLINE for CRV data format fragments. Did push in Mu2e_artdaq which has artdaq fragments. CRV digis to binaries, and art fragments to CRV digis. Eric says Giani is working on unpacking DAQ fragments by ROC - it should significantly reduced the data size in software.

Ron --